Prove that

d/dx (Yx) = 1/h (Yx+h – Yx–h) –1/2h (Yx+2h – Yx–2h) +1/3h (Yx+3h – Yx–3h) – ......... ...



1
Expert's answer
2021-01-10T17:23:01-0500

"\frac{y(x+h)-y(x-h)}{h}-\frac{y(x+2h)-y(x-2h)}{2h}+\frac{y(x+3h)-y(x-3h)}{3h}-...="


"=2y'-y'+2y'\/3-2y'\/4+2y'\/5-...=2y'(1-1\/2+1\/3-1\/4+1\/5)="


"=2y'ln2=1.386y'"